Natasha123:

There were five football games at a school . The number of hot dogs sold at the first four games is shown 523,604,544,596 the mean number of hot dogs sold at all five games was 573. How many hot dogs were sold at the fifth football game

AngeI:

Hello welcome to QuestionCove O: The mean of these numbers means the average (so you would add them all up and divide by the amount of different numbers you added together) If the mean of 5 games is 573, that means you would multiply 573 by 5 in order to get the total number of hotdogs sold in all 5 games. 573*5=2865 So next we would added together these 4 numbers, 523,604,544,596, in order to find out how many was sold in 4 games. 523+604+544+596=2267 You would then subtract 2865 by 2267 to get the number of hotdogs sold in game 5 2865-2267= 598 hot dogs